# Loop Series Expansions for Tensor Networks

### v1

@@ Appendix B Loop Excitations in the Thermodynamic Limit
In any finite tensor network it is possible, at least in principle, to enumerate over the different possible loop configurations, organize them by degree, and then evaluate the terms out to some desired degree. In contrast, in the thermodynamic limit, even counting the density of terms of some fixed degree may be an intractable combinatorial problem. To this end we now discuss two different approximations which can be applied to use the loop expansion to calculate the free energy density $f$, as defined in Eq. 10, of a homogeneous tensor network on a lattice in the thermodynamic limit (e.g. an iPEPS).
@@ B.1 Single-excitation approximation
The first method that we propose to estimate the contribution of loop excitations to the free energy density involves counting only the configurations that involve only a single excitation on top of the BP vacuum. Let $\mathcal{T}_{N}$ be a homogeneous tensor network on a square lattice of $N$ sites which is assumed to evaluate to some scalar $\mathcal{Z}_{N}$. We focus on the task of evaluating the free energy density $f_{N}$ defined
$f_{N}\equiv-\frac{\log(\mathcal{Z}_{N})}{N}.$ (16)
Let us assume that we have already found the BP fixed point of $\mathcal{T}_{N}$, and that the network has been normalized such that the BP vacuum is unity, $\mathcal{Z}_{\textrm{BP}}=1$ (or equivalently that the Bethe free energy is zero).
We shall compute the contributions from configurations that have a single excitation, with all other tensor edges set in their BP ground state. For simplicity, here we only consider the corrections arising from the $\delta_{4}$ excitations (occupying a single square on the lattice) although the method we describe is easily extendable to include arbitrary loop excitations. On the lattice of $N$ sites there are $N$ distinct locations where the $\delta_{4}$ excitation could be placed, see Fig. 7(a), with each configuration having a weight of $W(\delta_{4})$. It follows that the scalar $\mathcal{Z}_{N}$ of the network can be approximated as
$\mathcal{Z}_{N}\approx 1+NW(\delta_{4}),$ (17)
thus the free energy density can be further approximated as
$f_{N}$ $\approx\frac{-\log(1+NW(\delta_{4}))}{N}$
$\approx-W(\delta_{4}).$ (18)
To get to the second line of working in Eq. 18 we have used $\log(1+\epsilon)\approx\epsilon$ which requires $1/N\gg W(\delta_{4})$ to be valid. That the single-excitation approximation is only valid for small $N$ makes sense; for larger $N$ lattices the contributions from configurations with double-excitations (or higher) will dominate as their multiplicity scales at least as $N^{2}$. However, given that the result from Eq. 18 is independent of $N$, it may still be used as an approximation to the free energy density $f$ in the thermodynamic limit. Another way of interpreting this result is that we use the single-excitation approximation to calculate the free energy density on only a small patch of the otherwise infinite system. We can easily generalize Eq. 18 to include arbitrary excitations $\delta_{x}$,
@@
Figure 7: (a) A tensor network on a square lattice of $N=16$ sites is approximated via the single-excitation approximation for the $\delta_{4}$ loop excitations. The tensor indices in each expansion term, except those which are part of the $\delta_{4}$ excitation, are set in the BP ground state. (b) The multi-excitation approximation is used to expand the tensor network. In each expansion term (except for the BP vacuum) only the tensor indices directly connected to the $\delta_{4}$ excitation are set in the BP ground state. (c-d) Examples of terms which are double-counted in the multi-excitation approximation.
B.2 Multi-excitation approximation
The second method that we propose to estimate the contribution of loop excitations to the free energy density attempts to directly include contributions from configurations that have multiple excitations. For simplicity we again focus on the $\delta_{4}$ excitations, noting that the derivation is easily extended to include arbitrary excitations. We now resolve the tensor network $\mathcal{T}_{N}$ as the BP contribution $\mathcal{Z}_{\textrm{BP}}$ plus $N$ distinct contributions arising from the placement of a $\delta_{4}$ excitation on the lattice as shown Fig. 7(b). However, here in each contribution we freeze only edges connecting the excitation to their BP ground while leaving the other edges in the network free. It follows that each of these contributions has total weight $W(\delta_{4})\tilde{\mathcal{Z}}_{N-4}$, where $\tilde{\mathcal{Z}}_{N-4}$ represents the scalar from contracting the remaining tensor $\tilde{\mathcal{T}}_{N-4}$ network on $(N-4)$ sites. Thus the scalar $\mathcal{Z}_{N}$ from the network contraction is approximated as
$\mathcal{Z}_{N}\approx 1+NW(\delta_{4})\tilde{\mathcal{Z}}_{N-4}.$ (20)
Before continuing our derivation it is informative to reflect on the advantages and disadvantages of the proposed approximation. A key advantage of this proposal is that it implicitly includes multi-excitations; when resolving a contribution containing a single $W(\delta_{4})$ excitation the remaining network $\tilde{\mathcal{T}}_{N-4}$ could subsequently be resolved into a further sum of excitations. However, a significant disadvantage is that this approximation counts certain configurations multiple times, see Fig. 7(c-d) for examples. While it could be possible to use a more sophisticated strategy that subtracts out contributions from multiply-counted configurations, here we simply ignore them as in practice they seem to have minimal impact on accuracy.
In order to make headway with Eq. 20 another approximation is needed; we assume that the remaining network $\tilde{\mathcal{T}}_{N-4}$ has the same free energy density $f_{N}$ as the original network ${\mathcal{T}}_{N}$. In terms of the respective contraction scalars, $\tilde{\mathcal{Z}}_{N-4}$ and ${\mathcal{Z}}_{N}$, this implies
$\tilde{\mathcal{Z}}_{N-4}\approx e^{4f_{N}}{\mathcal{Z}}_{N}.$ (21)
Note that this approximation for $\tilde{\mathcal{Z}}_{N-4}$ ignores the boundary effects (i.e. resulting from fixing some of the network edges in their BP ground). Substitution of Eq. 21 into Eq. 20 gives
$\mathcal{Z}_{N}\approx 1+\mathcal{Z}_{N}\left(e^{4f_{N}}NW(\delta_{4})\right),$ (22)
which can be rearranged to give
$\mathcal{Z}_{N}\approx\left(1-e^{4f_{N}}NW(\delta_{4})\right)^{-1}.$ (23)
Finally, the free energy density $f_{N}$ can be expressed as
$f_{N}$ $\approx\frac{\log(1-e^{4f_{N}}NW(\delta_{4}))}{N}$
$\approx-e^{4f_{N}}W(\delta_{4}),$ (24)
@@ B.2 Multi-excitation approximation
with $L(\delta_{x})$ as the number of locations per lattice site that the excitation can be placed and $S(\delta_{x})$ as the number of lattice sites that the excitation occupies. Notice that Eq. 25 is identical to Eq. 19 apart from the addition of the extra terms $\exp{(S(\delta_{x})f)}$, which we refer to as bulk suppression factors as they act to suppress excitations that occupy larger regions of the network. The appearance of these factors in the multi-excitation derivation makes intuitive sense: excitations that occupy larger regions leave less ‘room’ for other excitations to occur and are penalised accordingly (whereas this consideration is irrelevant if only single excitations are allowed).
@@
Figure 8: Results comparing the accuracy of (a) the single-excitation approximation and (b) the multi-excitation approximation for computing the free energy density $f$ of the hexagonal AKLT model when including all excitations $\delta_{x}$ up to degree $x$. While both approximations give similar results for the $\delta_{6}$ excitation, the single-excitation approximation fails to converge to the exact result as higher-degree excitations are included.
B.3 Comparison
We have presented two approaches, each based on different approximations, for using the loop expansion to estimate the free energy density on an infinite, homogeneous tensor network. A comparison of these approaches for the hexagonal lattice AKLT model is shown in Fig. 8. While both approximations show similar accuracy when the expansion includes only the $\delta_{6}$ excitation, the accuracy of single-excitation approximation shows minimal improvement as higher-degree terms are included in the expansion. In contrast, the multi-excitation approximation appears to converge to the exact result as the expansion is taken to include higher-degree terms. Similar results were also observed for the AKLT models on square and kagome lattices, and for randomly initialized PEPS tensor networks. These results demonstrate that the bulk suppression factors $\exp{(S(\delta_{x})f)}$, which appear only in Eq. 25 derived using the multi-excitation approximation, are vital to converge the loop expansions to high accuracy.
@@ Appendix C Kagome and Square Lattices
In this appendix we provide some of the details of the benchmark results provided for iPEPS on kagome and square lattices. An iPEPS for the kagome lattice is shown in Fig. 9(a). Although we could perform the series expansion on the (square norm) of this network directly, we instead found it preferable to perform some preliminary restructuring. The reasoning for this was to remove the triangular loops from the network which are otherwise problematic since (i) the weight of the excitations around the triangular loops are likely to be large (since the loops are short) and (ii) they generate a multitude of low-degree terms that would need to be included in the expansion. The restructuring is depicted in Fig. 9(b), where each (5-index) PEPS tensor is first decomposed as a product of three 3-index tensors, then the tensors surrounding each triangle of the kagome lattice are subsequently contracted into a single tensor (located at the center of each kagome lattice triangle). These manipulations result in a PEPS with a hexagonal lattice geometry but where the physical indices are located on tensors set on the edges of the hexagonal lattice as shown in Fig. 9(c). Upon forming the square norm $\langle\psi|\psi\rangle$ a purely hexagonal lattice tensor network is obtained, as shown in Fig. 9(d), which can be treated with the same expansion as described for the hexagonal lattice PEPS. Note that the manipulations performed in Fig. 9(b) may require an increase in PEPS bond dimension from $m\to m^{2}$, although in many cases the new bond dimension could be truncated without incurring significant error. This example of manipulating the kagome lattice into a hexagonal lattice highlights an interesting point: that the accuracy of the series expansion can depend on the particulars of the tensor network representation of a quantum state $|\psi\rangle$, rather than solely on the properties of $|\psi\rangle$ itself (e.g. entanglement and correlations).

### v2

@@ Appendix B Loop Excitations in the Thermodynamic Limit
In any finite tensor network it is possible, at least in principle, to enumerate over the different possible loop configurations, organize them by degree, and then evaluate the terms out to some desired degree. In contrast, in the thermodynamic limit, even counting the density of terms of some fixed degree may be an intractable combinatorial problem. To this end, we now discuss three different methods to approximate the loop corrections to the free energy density $f$, as defined in Eq. 10, of a homogeneous tensor network on a lattice in the thermodynamic limit (e.g. an iPEPS).
@@ B.1 Single-excitation approximation
The first method that we propose to estimate the contribution of loop excitations to the free energy density involves counting only the configurations that involve only a single excitation on top of the BP vacuum. Let $\mathcal{T}_{M}$ be a homogeneous tensor network on a hexagonal lattice of $M$ plaquettes that, when contracted, evaluates to some scalar $\mathcal{Z}_{M}$. We begin by assuming that we have already found the BP fixed point of the network, and that the network has been normalized such that the BP vacuum is unity, $\mathcal{Z}_{\textrm{BP}}=1$ (or equivalently that the Bethe free energy is zero).
We now consider evaluating the corrections arising from the $\delta_{6}$ excitations, see Fig. 7(a), although our derivation is easily extendable to include arbitrary loop excitations. Given that there are $M$ distinct locations that the $\delta_{6}$ excitation could occupy, with each configuration having a weight of $W(\delta_{6})$, it follows that the scalar $\mathcal{Z}_{M}$ of the network is
$\mathcal{Z}_{M}\approx 1+MW(\delta_{6}).$ (16)
The (per-site) free energy density can thus be approximated as
$f$ $\approx\frac{-\log\left(1+MW(\delta_{6})\right)}{2M}$
$\approx-\frac{1}{2}W(\delta_{6}).$ (17)
To get to the second line of working in Eq. 17 we have used $\log(1+\epsilon)\approx\epsilon$ which requires $1/M\gg W(\delta_{6})$ to be valid. That the single-excitation approximation is only valid for small $M$ makes sense; for larger $M$ lattices the contributions from configurations with double-excitations (or higher) will dominate as their multiplicity scales at least as $M^{2}$. However, given that the result from Eq. 17 is independent of $M$, it may still be used as an approximation to the free energy density $f$ in the thermodynamic limit. Another way of interpreting this derivation is that we are using the single-excitation approximation to calculate the free energy density on only a small patch of the otherwise infinite system. We can easily generalize Eq. 17 to include arbitrary excitations $\delta_{x}$,
@@
Figure 7: (a) A hexagonal lattice tensor network is approximated via the single-excitation approximation for the $\delta_{6}$ loop excitations. The tensor indices in each expansion term, except those which are part of the $\delta_{6}$ excitation, are set in the BP ground state. (b) The self-consistent completion is used to expand the tensor network for the $\delta_{6}$ loop excitations. In each expansion term only the tensor indices directly connected to the $\delta_{6}$ excitation are set in the BP ground state.
B.2 Self-consistent completion
The second method that we propose to estimate the contribution of loop excitations uses a self-consistent completion. For simplicity, we again focus on the $\delta_{6}$ excitations, noting that the derivation is easily extended to include arbitrary excitations. We now resolve the hexagonal network of $M$ plaquettes as the BP contribution $\mathcal{Z}_{\textrm{BP}}$ plus $M$ distinct contributions arising from the placement of a $\delta_{6}$ excitation on the lattice as shown Fig. 7(b). However, in each contribution we freeze only edges connecting the excitation to their BP ground while leaving the other edges in the network free. It follows that each of these contributions has total weight $W(\delta_{6})\tilde{\mathcal{Z}}_{M-7}$, where $\tilde{\mathcal{Z}}_{M-7}$ represents the scalar from contracting the remaining tensor network, which now consists of $(M-7)$ complete hexagonal plaquettes. Thus the scalar $\mathcal{Z}_{M}$ from the network contraction is approximated as
$\mathcal{Z}_{M}\approx 1+MW(\delta_{6})\tilde{\mathcal{Z}}_{M-7}.$ (19)
In order to make headway with Eq. 19 another approximation is needed; we now assume that $\tilde{\mathcal{Z}}_{M-7}$ and ${\mathcal{Z}}_{M}$ have the same free energy density $f$, which implies
$\tilde{\mathcal{Z}}_{M-7}\approx e^{7f}{\mathcal{Z}}_{M}.$ (20)
Note that this approximation for $\tilde{\mathcal{Z}}_{M-7}$ ignores boundary effects (i.e. resulting from fixing some of the network edges in their BP ground). Substitution of Eq. 20 into Eq. 19 gives
$\mathcal{Z}_{M}\approx 1+\mathcal{Z}_{M}\left(e^{7f}MW(\delta_{6})\right),$ (21)
which can be rearranged as
$\mathcal{Z}_{M}\approx\left(1-e^{7f}MW(\delta_{6})\right)^{-1}.$ (22)
Finally, the (per-site) free energy density $f$ can be expressed as
$f$ $\approx\frac{\log(1-e^{7f}MW(\delta_{6}))}{2M}$
$\approx-\frac{1}{2}W(\delta_{6})e^{7f},$ (23)
@@ B.2 Self-consistent completion
with $L(\delta_{x})$ as the number of locations per lattice site that the excitation can be placed and $S(\delta_{x})$ as the number of hexagonal plaquettes that each excitation occupies. Notice that Eq. 24 is identical to Eq. 18 apart from the addition of the suppression factors $\exp{(S(\delta_{x})f)}$, which serve to approximate the contributions arising from configurations containing multiple excitations.
@@
B.3 Multi-excitation counting
A third strategy to estimate the contribution of loop excitations is simply to explicitly count the products of distinct excitations  [^1] I. Lukin and A. Sotnikov, private communication.. Again we consider a network on a hexagonal lattice of $M$ plaquettes that evaluates to some scalar $\mathcal{Z}_{M}$. Here the loop corrections, when expanded out to the degree-10 terms, remain identical to those from the single-excitation approximation in Eq. 18,
$\mathcal{Z}_{M}=MW(\delta_{6})+3NW(\delta_{10}).$ (25)
However, the corrections from degree-12 excitations, $(\Delta\mathcal{Z})_{12}$, now include the $\delta_{12}$ excitations in addition to the product of two degree-6 excitations $\delta_{6}$,
$(\Delta\mathcal{Z})_{12}=2MW(\delta_{12})+\frac{M(M-7)}{2}W(\delta_{6})^{2},$ (26)
where the factor $\frac{M(M-7)}{2}$ counts the number of distinct locations that the two degree-6 excitations can be placed. The corrections from the degree-14 excitations remain unchanged,
$(\Delta\mathcal{Z})_{14}=12NW(\delta_{14}),$ (27)
but the degree-16 excitations additionally include the product of $\delta_{6}$ and $\delta_{10}$ excitations,
$(\Delta\mathcal{Z})_{16}=18MW(\delta_{16})+3M(M-10)W(\delta_{6})W(\delta_{10}).$ (28)
This methodology, which explicitly counts product excitations, can easily be extended out to arbitrary degree, e.g. where the degree-18 expansion would also include products of three $\delta_{6}$ excitations.
Taking the logarithm of $\mathcal{Z}_{M}$, the (per-site) free energy density $f$ is given as
$f=$ $-\frac{1}{2}W(\delta_{6})-\frac{3}{2}W(\delta_{10})-W(\delta_{12})+$
$\frac{7}{4}W(\delta_{6})^{2}-6W(\delta_{14})-9W(\delta_{16})+$
$15W(\delta_{6})W(\delta_{10})+\ldots,$ (29)
where it is noted that the terms of order $M^{2}$ canceled out in the logarithm expansion.
B.4 Comparison
We have presented three approaches for using the loop expansion to estimate the free energy density on an infinite, homogeneous tensor network. A numerical comparison of these approaches is shown in Fig. 8 for loop corrections (expanded out to degree-16) in the hexagonal lattice AKLT model. While the expansion using the single-excitation approximation saturates at a finite accuracy, both the self-consistent completion and the multi-excitation approach appear to converge to the exact result as the degree of the expansion is increased. Similar results were also observed for the AKLT models on square and kagome lattices, and for randomly initialized PEPS tensor networks.
While both the self-consistent completion and the multi-excitation counting are seen to give comparable accuracy, the numerical results of the main text were derived using the former approach. A reason for this is that the self-consistent completion is easier to apply to the case of computing transfer and density matrices, given that the same suppression factors $\exp{(S(\delta_{x})f)}$ can be used as for the free energy density calculation.
Figure 8: Results comparing the accuracy of (a) the single-excitation approximation, (b) the self-consistent completion, and (c) the multi-excitation counting strategies for computing the free energy density $f$ of the hexagonal AKLT model when including all excitations $\delta_{x}$ up to degree $x$. While the single-excitation approximation saturates at a finite accuracy, the other strategies appear to converge to the exact result in the limit of a large degree expansion.
@@ Appendix C Kagome and Square Lattices
In this appendix we provide some of the details of the benchmark results provided for iPEPS on kagome and square lattices. An iPEPS for the kagome lattice is shown in Fig. 9(a). Although we could perform the series expansion on the (square norm) of this network directly, we instead found it preferable to perform some preliminary restructuring. The reasoning for this was to remove the triangular loops from the network which are otherwise problematic since (i) the weight of the excitations around the triangular loops are likely to be large (since the loops are short) and (ii) they generate a multitude of low-degree terms that would need to be included in the expansion. The restructuring is depicted in Fig. 9(b), where each (5-index) PEPS tensor is first decomposed as a product of three 3-index tensors, then the tensors surrounding each triangle of the kagome lattice are subsequently contracted into a single tensor (located at the center of each kagome lattice triangle). These manipulations result in a PEPS with a hexagonal lattice geometry but where the physical indices are located on tensors set on the edges of the hexagonal lattice as shown in Fig. 9(c). Upon forming the square norm $\langle\psi|\psi\rangle$ a purely hexagonal lattice tensor network is obtained, as shown in Fig. 9(d), which can be treated with the same expansion as described for the hexagonal lattice PEPS. Note that the manipulations performed in Fig. 9(b) may require an increase in PEPS bond dimension from $m\to m^{2}$, although in many cases the new bond dimension could be truncated without incurring significant error. This example of manipulating the kagome lattice into a hexagonal lattice highlights an interesting point: the accuracy of the series expansion can depend on the particulars of the tensor network representation of a quantum state $|\psi\rangle$, rather than solely on the properties of $|\psi\rangle$ itself (e.g. entanglement and correlations).
@@
Figure 11: (a) Definition of a corner double line (CDL) tensor. (b) A transfer matrix $T$ formed by opening a single index from an infinite network of CDL tensors. (c) The transfer matrix $T$ is exactly computed from the BP vacuum plus the single and double loop corrections. (d) Definition tensor $Y$. (e) Definition of isometry $W$. (f-g) A boundary MPS, formed from copies of $Y$, is seen to be an exact fixed point of the boundary MPS contraction when implemented in conjunction with isometries $W$.
Appendix D Contraction Efficiency for PEPS versus MPS-based Approaches
In the main text we demonstrated the use of the loop expansion to approximately contract PEPS tensor networks, although we did not attempt to quantify the efficiency (in terms of accuracy versus computational cost) versus established PEPS contraction methods such as boundary MPS or corner transfer matrix renormalization group (CTMRG). One reason for this was that a proper comparison would require a thorough and dedicated investigation as the relative accuracy is likely highly variable on the test model under consideration, and also that the cost scaling of MPS-based methods can depend on the finer details of their implementation. Given that our aim was to present the loop expansion as a general method to approximate arbitrary tensor network contractions, with the application to contract PEPS only as a single example, we felt such a comparison was beyond the scope of the present work. None-the-less, in this appendix we present an example that demonstrates that there exist $2D$ networks that are demonstrably more efficient to contract with a loop expansion than with previous MPS-based approaches.
Consider a homogeneous square-lattice tensor network of bond dimension $D$ formed from corner double-line (CDL) tensors, themselves composed of corner-sharing entangled pairs mediated by matrices $\sigma$ of dimension $\sqrt{D}$ as shown in Fig. 11(a). We assume that the $\sigma$ matrices are diagonal and contain an exponentially decaying weights,
$\sigma_{ij}=\delta_{ij}\exp(-\alpha i)$ (30)
with $\alpha$ a positive constant. We focus on the task of computing the transfer matrix $T$ formed by cutting open a single index from the infinite network as shown in Fig. 11(b). We begin by computing $T$ using the loop series expansion, where computing the series expansion out to the double-plaquette term, see Fig. 11(c), is found to give the exact result. The cost of evaluating the double-plaquette correction is $O(D^{4})$, although this cost reduces to $O(D^{3})$ if only the single-plaquette loop corrections are considered, which may still provide a good approximation to $T$.
Figure 12: Leading order computational cost versus accuracy for the error in the transfer matrix, $||T-T_{\textrm{exact}}||$, from a network of CDL tensors of dimension $D=36$ for (a) $\alpha=0.1$ and (b) $\alpha=0.4$ as defined in Eq. 30. Compared are bond dimensions $m=1,2,\ldots,6$ for the boundary MPS versus the loop expansion with either single or double loop corrections. The results demonstrate that the loop expansion is the more efficient approach: it produces higher accuracy for lower computational cost.
We now consider calculating $T$ through a boundary MPS approachPEPS5 . Alternatively, a corner transfer matrix approachPEPS6 could be used, although this is found to be equivalent in efficiency for the network under consideration. It can be seen that an MPS formed from tensors $Y$ as defined in Fig. 11(d), in conjunction with isometries $W$ as defined in Fig. 11(d), produces an exact fixed-point boundary MPS as shown in Figs. 11(f-g). The transfer matrix $T$ can then be evaluated from (two copies of) the boundary MPS using standard methodology. If the bond dimension $m$ of the boundary MPS is fixed at $m=\sqrt{D}$ then the contraction for $T$ is exact, although accurate approximations to $T$ can be obtained with smaller $m$. The leading order cost of contracting for $T$ using the boundary MPS approach is known to scale as $O(D^{4}m^{2})$.
In Fig. 12 we compare the efficiency of the loop expansion versus the boundary MPS approach (in terms of the error $||T_{\textrm{e}xact}-T_{\textrm{approx}}||$ versus the leading order computational cost) for several values of $\alpha$ in Eq. 30. It can be seen that the loop expansion is the more efficient approach for contracting the CDL network: it produces more accurate results with less computational effort over the values of $\alpha$ considered. Although the CDL tensor network only has a simple entanglement structure, this example none-the-less demonstrates that the loop expansion may offer a more efficient contraction than traditional MPS-based approaches for contraction of certain $2D$ tensor networks.